OpenStudy (anonymous):

integral 1 to 9 (x-1)/ sqrt x dx

OpenStudy (jamesj):

rmalik, this is the 5th or 6th integration problem you've posted. This one is solved in the same way as all the others.

myininaya (myininaya):

write as two fractions

OpenStudy (jamesj):

So I suggest you go back first and make sure you understand the solutions to all those other problems. If you do, you'll have no problem with one.

OpenStudy (anonymous):

but its difficult...how..i know sqrt x is x^1/2

OpenStudy (jamesj):

Yes, so \[\frac{x-1}{\sqrt{x}} = x^{1/2} - x^{-1/2}\]

OpenStudy (jamesj):

So now, if you understand the other problems, you should know how to integrate this.

OpenStudy (anonymous):

but what happens with the numerator?

OpenStudy (jamesj):

\[\frac{x-1}{\sqrt{x}} = \frac{x}{\sqrt{x}} - \frac{1}{\sqrt{x}}\]

OpenStudy (jamesj):

Now, \[\sqrt{x} = x^{1/2}\] and \frac{1}{\sqrt{x}} = \frac{1}{x^{1/2}} = x^{-1/2}

OpenStudy (jamesj):

\[\frac{1}{\sqrt{x}} = \frac{1}{x^{1/2}} = x^{-1/2}\]

OpenStudy (jamesj):

So that is how I changed the function you are integrating into the form a few posts above

OpenStudy (jamesj):

Now ... let me ask you this. In general, what is \[\int\limits x^p \ dx\] where p is not -1 ?

OpenStudy (jamesj):

what is this indefinite integral?

OpenStudy (jamesj):

all of your exercises have depended on this integral

OpenStudy (anonymous):

x^2/2

OpenStudy (jamesj):

That's right for p = 1. But what about p in general

OpenStudy (anonymous):

i dont know

OpenStudy (jamesj):

Right, so to get to that answer, can you tell me what is the derivative of x^n? \[\frac{d \ }{dx} x^n\]

OpenStudy (anonymous):

nx

OpenStudy (jamesj):

Not quite.

OpenStudy (anonymous):

or would it be x^n+1

OpenStudy (jamesj):

not quite

OpenStudy (anonymous):

if you dont want to help me dont bother

OpenStudy (jamesj):

\[\frac{d \ }{dx} x^n = nx^{n-1}\]

OpenStudy (jamesj):

I am helping you. Be patient. Do you recognize this formula?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Right. Now, integration is "anti-differentiation", so what is the integral of x^{n-1} \[\int\limits x^{n-1} dx\]

OpenStudy (jamesj):

i.e., what function when differentiated gives x^(n-1) ?

OpenStudy (anonymous):

but im not understanding the problem i asked...since its x^1/2-x^-1/2...then i solve the integral which would be 1^1/2-1^-1/2-9^1/2-9^-1/2

OpenStudy (jamesj):

In general, for p not equal to -1 \[\int\limits x^p \ dx = \frac{1}{p+1} x^{p+1} + C\]

OpenStudy (jamesj):

And that is because \[\frac{d \ }{dx} x^{p+1} = (p+1) x^p\]

OpenStudy (jamesj):

So now, with that formula you can calculate \[\int\limits (x^{1/2} - x^{-1/2}) \ dx \]

OpenStudy (anonymous):

yes like i showed above is that how i would do it?

OpenStudy (jamesj):

No, what you wrote above was wrong. That's why I took back into the basics here so you would understand why.

OpenStudy (jamesj):

Using the integration formula above for x^p, what is \[\int\limits x^{1/2} \ dx\] ?

OpenStudy (anonymous):

3x^2/3/2

OpenStudy (jamesj):

p = 1/2 so the integral is \[\frac{1}{1 + 1/2} x^{1 + 1/2} = \frac{1}{3/2} x^{3/2} = \frac{2}{3} x^{3/2}\]

OpenStudy (anonymous):

o i had it the wrong way

OpenStudy (jamesj):

What now is \[\int\limits x^{-1/2} \ dx\]

OpenStudy (anonymous):

hold on i will write on paper first

OpenStudy (anonymous):

it would be 2x^1/2?

OpenStudy (jamesj):

Yes.

OpenStudy (anonymous):

ok cool now please help with next steps i think i get it now

OpenStudy (jamesj):

So now you know how to do \[\int\limits_1^9 \frac{x -1}{\sqrt{x}} \ dx\]

OpenStudy (anonymous):

wait wait wait....so heres what i am doing....

OpenStudy (anonymous):

im sorry i dont

OpenStudy (jamesj):

what is the indefinite integral? \[\int\limits \frac{x-1}{\sqrt{x}} dx\] ?

OpenStudy (anonymous):

2x^1/2?

OpenStudy (anonymous):

so i would just plug in the intrgral og 0 to 9?

OpenStudy (jamesj):

If that were right, then the derivative of (2x^(1/2)) would be the function we're trying to integrate. But it's not. The derivative of (2x^(1/2)) is just x^(-1/2)

OpenStudy (jamesj):

What is the function we're trying to integrate?

OpenStudy (anonymous):

i really dont understand and im serious

OpenStudy (jamesj):

We are trying to integrate the function \[\frac{x-1}{\sqrt{x}}\]

OpenStudy (jamesj):

Now, as it stands, that's not a function we can deal with. So we figure out how to write in terms of things we understand.

OpenStudy (anonymous):

you said its = to x^1/2=x^-1/2

OpenStudy (jamesj):

\[\frac{x-1}{\sqrt{x}} = \frac{x-1}{x^{1/2}} = \frac{x}{x^{1/2}} - \frac{1}{x^{1/2}}\]

OpenStudy (jamesj):

And we can simplify that further \[\frac{x}{x^{1/2}} - \frac{1}{x^{1/2}} = x^{1-1/2} - x^{-1/2} = x^{1/2} - x^{-1/2}\]

OpenStudy (jamesj):

Do understand everything I just did?

OpenStudy (anonymous):

so it becomes x^1/2-x^-1/2..dx which result to become (2x^3/2)/3-2x^1/2

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

then i compute the intergral??

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

FINALLY!!!!!! haha

OpenStudy (jamesj):

Tell me when you have the answer.

OpenStudy (anonymous):

so it would be...well no answer yet...im up to (2^3/2)/3-(2^1/2)-(54/3)-6..im not sure if that's right

OpenStudy (jamesj):

Ok. Let's make sure agree on the indefinite integral \[\int\limits x^{1/2} - x^{-1/2} \ dx = \frac{2}{3} x^{3/2} - 2 x^{1/2} \ ( \ + C \ )\]

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

So now if you want the definite integral from x = 1 to 9, substitute x=9 of this expression minus x = 1 of this expression. What do you get when you substitute x = 1?

OpenStudy (anonymous):

(2(1)^3/2)/3-2(1)^1/2= 2/3 -2?

OpenStudy (jamesj):

yes ... which is ....?

OpenStudy (anonymous):

-4/3?

OpenStudy (jamesj):

of course.

OpenStudy (jamesj):

Now, what is the integral evaluated at x = 9?

OpenStudy (anonymous):

well i got 54/3-6

OpenStudy (anonymous):

(54/3)-6?

OpenStudy (jamesj):

which simplifies to ...

OpenStudy (anonymous):

35/3?

OpenStudy (anonymous):

which is 40/3?

OpenStudy (jamesj):

No ... 54 = 9 x 6 = 18 x 3 Hence 54/3 = 18

OpenStudy (jamesj):

Right. The definite integral is equal to 40/3.

OpenStudy (jamesj):

That is the final answer.

OpenStudy (anonymous):

ok wonderfull thanks sooooooooooooooooooooooo much!!

OpenStudy (jamesj):

So now, do yourself a favor. Take a blank piece of paper and work the entire problem again from the beginning. Then take another piece of blank piece of paper and work the entire problem without looking at the first piece or our discussion here. When you can do that, then you really understand this problem.

OpenStudy (anonymous):

ok great help thanks agian

OpenStudy (jamesj):

Also, go back and look at the other questions you've asked and make sure you understand the solutions.

OpenStudy (jamesj):

ok ... good luck.